# Med3DVLM: An Efficient Vision-Language Model for 3D Medical Image Analysis

### v2

@@ 1 Introduction
Medical image analysis plays a crucial role in diagnosing and treating diseases such as cancer, cardiovascular conditions, and neurological disorders. However, existing models are often task-specific, lack adaptability to new tasks, and do not support real-time user interactions. Vision-language models (VLMs), such as Contrastive Language-Image Pre-Training (CLIP) [28] and Large Language and Vision Assistant (LLaVA) [20], offer greater versatility in dynamic clinical settings by aligning medical images with textual reports. CLIP leverages contrastive learning on image-text pairs to enable zero-shot image classification and image-text retrieval. LLaVA extends CLIP by integrating its visual encoder with a large language model (LLM), facilitating interactive tasks such as report generation and visual question answering (VQA). Despite their success in analyzing 2D medical images, such as chest X-rays [8] and pathology slides [22], their application to 3D imaging remains limited.
@@
Several VLMs have attempted to bridge this gap, but each has limitations in the 3D setting. For example, PMC-CLIP [18], trained on large-scale biomedical literature images, is restricted to 2D inputs, leading to poor performance on 3D image understanding. RadFM [39] unifies 2D and 3D data but is primarily optimized for text generation tasks, such as VQA, and struggles with broader image-text understanding. More recently, M3D-LaMed [2] was introduced as a generalist multi-modal model for 3D medical image analysis. By combining a CLIP-pretrained 3D visual encoder with a 3D spatial pooling perceiver, M3D-LaMed enables direct reasoning on 3D scans and achieves state-of-the-art performance across multiple benchmarks, including image-text retrieval, report generation, open-ended and closed-ended VQA, as well as 3D segmentation and localization tasks.
@@
    Efficient 3D Feature Encoding. We integrate DCFormer [1], a 3D image encoder that efficiently captures volumetric features by decomposing 3D convolutions into three parallel 1D convolutions along the depth, height, and width axes. This approach reduces computational complexity and enables richer, more scalable feature representations for 3D image volumes.
@@
    Improved Vision-Language Alignment. We adopt SigLIP [45], a sigmoid-based language-image pretraining scheme, to improve image-text alignment. Unlike CLIP’s softmax-based contrastive loss, which relies on distinguishing positive pairs from a large batch of negative samples, SigLIP uses a pairwise sigmoid loss that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, making training more stable and less sensitive to batch size.
@@
    Multi-Scale Multi-Modal Projector. We introduce a novel projector based on MLP-Mixer [30] to fuse image and text embeddings effectively. Utilizing a low-high-hybrid design, it blends detailed low-level and abstract high-level features from the image encoder with LLM text embeddings using stacked MLP layers. Inspired by MLP-Mixer’s ability to mix spatial and feature data, this dual-stream approach captures richer cross-modal interactions than simple linear projection, improving the LLM’s decoding accuracy.
@@ 2.1 Medical Vision-Language Models
Medical VLMs utilize multi-modal learning to enhance tasks such as disease diagnosis, report generation, and medical VQA through improved integration and understanding of medical images and text. One of the earliest efforts, ConVIRT [47], applied contrastive learning to learn visual representations from paried image-text data, outperforming ImageNet pretraining in medical image classification and zero-shot retrieval tasks. Similarly, BioViL [4] leveraged large-scale biomedical datasets to refine multi-modal representations, significantly outperforming previous supervised methods. Further extending these capabilities, MedCLIP [36] replaced the InfoNCE loss with semantic matching loss based on medical knowledge, demonstrating its ability to learn generalized representations with limited data.
Recent efforts have aimed to develop VLMs for 3D medical imaging. RadFM [39] introduced a generalist VLM trained on massive datasets that consisted of both 2D and 3D medical images and associated radiology reports. It integrated contrastive learning and generative modeling across diverse imaging modalities (X-ray, CT, MRI) and textual reports, enabling a unified radiology representation. Extending this, M3D [2] introduced a large-scale dataset, M3D-Data, along with M3D-LaMed, a multi-modal vision-language model for 3D medical image analysis tasks such as image-text retrieval, report generation, VQA, and promotable segmentation. Its successor, E3D-GPT [15], employed a 3D multi-modal masked autoencoder framework to further enhance image-text retrieval, leading to improved performance in report generation and VQA. In parallel, researchers have also explored text-guided 3D medical image segmentation [48, 14, 40], where language prompts are used to localize anatomical structures or pathologies.
@@ 2.2 Radiology Report Generation
Automated report generation aims to produce descriptive, accurate, and clinically relevant reports from medical images. It improves diagnostic accuracy while reducing radiologists’ workload and supporting care in resource-limited settings. Early approaches primarily adopted an encoder-decoder architecture, where convolutional neural networks (CNNs) extracted image features, and long short-term memory (LSTM) networks generated text descriptions [41, 12, 44, 37]. However, these methods struggled with long-range dependencies, often produced repetitive text, and had limited capacity to capture complex medical semantics.
The transformer architecture [33] addressed key limitations of CNNs and LSTMs by using self-attention to capture global dependencies. Pretrained LLMs [31, 27] based on transformers are now adapted into medical VLMs for report generation. R2GenGPT [35] introduced an efficient visual alignment module to better integrate image features with LLM word embeddings, improving text coherence and clinical relevance. Med-Flamingo [23], based on OpenFlamingo-9B, was pretrained on interleaved medical image-report pairs, achieving superior performance in generating clinically useful responses. More recent methods, such as LLaVA-Med, were based on the LLaVA [20] framework, which combined a visual encoder with an LLM to generate detailed and coherent reports. CT2Rep [11] proposed an advanced 3D vision encoder to generate radiology reports for 3D medical imaging, specifically targeting chest CT volumes. Similarly, CT-CHAT [10] adapted the LLaVA framework for chest CT report generation, demonstrating the effectiveness of large-scale pretrained LLMs in 3D medical imaging.
@@ 2.3 Medical Visual Question Answering
A medical VQA system can answer natural language questions about medical images. This task can be closed-ended (with answers like yes/no or a choice from a fixed list) or open-ended (free-form text answers). Early work on medical VQA largely treated it as a classification problem, where models selected the correct answer from a predefined set of possible responses [24, 49, 9, 19]. These models typically encoded images (e.g., using CNNs) and text (e.g., using transformers), and then mapped the combined features to a predefined answer space. Although this approach worked well for simple questions, such as identifying an organ, it struggled with questions requiring detailed explanation or answers not included in the predefined list.
Nowadays, approaches to medical VQA increasingly adopt generative architectures, treating answer generation as a sequence prediction task, which enables open-ended responses. An early step in this direction was CGMVQA [29], which added a generative decoder branch alongside the traditional classifier. Recent VQA systems have fully transitioned to generative models [7]. These models (e.g., Med-PaLM M [32] and LLaVA-Med [16]) extend an LLM with visual inputs via visual instruction tuning [20], allowing the LLM to interpret medical images and answer questions in a conversational manner. Recently, a few works have explored VQA on 3D medical images [2, 10], where models can directly perceive entire 3D image volumes and generate context-aware responses, offering more accurate and holistic medical insights.
@@ 3.1 Dataset
This study used the publicly available M3D dataset [2], collected with informed consent and ethical approval by the original investigators. All data were de-identified, and no additional consent or approval was required for this secondary analysis. We used two subsets of the M3D dataset: M3D-Cap and M3D-VQA. M3D-Cap includes 120K image-text pairs, while M3D-VQA contains 662K instruction-response pairs. The dataset can support tasks such as image-text retrieval, radiology report generation, and VQA. All image volumes were resampled to a fixed size of 128x256×256. The dataset was divided into training (115k samples), validation (3k samples), and test (2k samples) sets.
@@ 3.2.1 Contrastive Pretraining
Following the CLIP framework [28], the vision and text encoders are trained jointly using a contrastive loss to align 3D images with their corresponding radiology reports. The goal is to maximize the similarity between embeddings of positive (matching) image-text pairs while minimizing it for negative (non-matching) pairs.
For the vision encoder, we utilize DCFormer [1], which efficiently extracts volumetric features using decomposed 3D convolutions. This approach mitigates computational challenges in high-resolution 3D image analysis, where traditional 3D CNNs and vision transformers struggle due to the cubic scaling of 3D CNNs with kernel size and the quadratic scaling of vision transformers with image size. DCFormer decomposes 3D convolutions into three 1D parallel depthwise convolutions:
@@
Such a decomposition strategy enables the processing of 3D volumes at a larger size of 128x256×256 in this study, preserving fine-grained spatial details for improved image-text alignment. DCFormer, short for DeComposed Former, derives its name from this decomposition strategy. The “Former” suffix is inspired by the MetaFormer [43] framework, which generalizes Transformer-like architectures by replacing attention mechanisms with alternative token-mixing operations.
In this work, we utilize DCFormer-small variant, which consists of a stem stage and four hiararchical stages with [2, 3, 6, 2] layers and [96, 192, 384, 768] channels [1]. This adaptation of DCFormer within the LLaVA [20] framework enables efficient processing of high-resolution 3D volumes while preserving spatial detail. By reducing computational complexity, it supports scalable 3D medical vision-language modeling under typical GPU constraints and addresses key limitations of 3D vision transformers and CNNs.
For the text encoder, we use ClinicalBERT [34], an adaptation of BERT [6] pretrained on clinical notes. Designed for clinical text, ClinicalBERT is well-suited for processing radiology reports, generating embeddings that effectively capture their semantic content.
Traditional softmax-based contrastive learning (e.g., CLIP) distinguishes positive pairs from a large batch of negative samples. While effective for large datasets, this approach struggles with smaller medical datasets, where meaningful negative samples are limited, and similar images may share semantic information. To address this, we adopt SigLIP [45], a sigmoid-based language-image pretraining scheme that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, resulting in more stable training and reduced sensitivity to batch size. Sigmoid loss is defined as:
$\mathcal{L}_{\text{SigLIP}}=-\frac{1}{|B|}\sum_{i=1}^{|B|}\sum_{j=1}^{|\mathcal{B}|}\log\underbrace{\frac{1}{1+e^{z_{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}_{\mathcal{L}_{ij}},$ (5)
where $z_{ij}$ is the label for the $i$-th image and $j$-th text pair, $z_{ij}=1$ for positive pairs and $z_{ij}=0$ otherwise. The sigmoid loss $\mathcal{L}_{ij}$ is computed for each image-text pair, with $t$ as the temperature parameter, $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ as the image and text embeddings, respectively, and $b$ as a bias term. The overall loss is averaged over all pairs in the batch.
@@ 3.2.2 Multi-modal Projector Pretraining
We propose an effective multi-modal projector for the interaction between visual and textual embeddings. In this pretraining stage, all model weights are frozen except for those in the projector. To effectively capture spatial and feature dependencies, we adopt the MLP-Mixer architecture [30], which alternates between a token-mixing MLP (spatial mixer) and a channel-mixing MLP.
@@
To further enhance multi-modal fusion, we propose a low-high hybrid Mixer-MLP architecture (Figure 2), inspired by Janus Pro [5]. Unlike M3D-LaMed [2], which produces large spatial outputs $(B,2048,768)$ that require lossy downsampling, DCFormer outputs semantically rich features $(B,32,768)$ from its final layer and spatially detailed features $(B,256,384)$ from its penultimate layer. Because of the compact spatial dimensions of DCFormer’s outputs, our model can adopt the low-high hybrid structure without exceeding the token length limit of the LLM. This is crucial, as excessively long image token sequences can crowd out text tokens, leading to truncated questions or reports during inference. To leverage both types of features, we utilize two parallel MLP-Mixer modules—one for each layer. After passing through $N$ Mixer layers, the resulting feature outputs are concatenated into a unified sequence of image tokens. In parallel, the input text is tokenized into a sequence of token IDs using the vocabulary of the LLM. This step follows the LLaVA [20] framework, which requires token embeddings for multimodal fusion. The resulting image and text tokens are then fused and passed to the LLM for joint reasoning.
@@ 3.2.3 VLM Fine-tuning
In this stage, we fine-tune the multi-modal projector and the LLM for the report generation and VQA. Instead of fine-tuning the entire LLM, we adopt Low-Rank Adaptation (LoRA) [13], which significantly reduces the number of trainable parameters by introducing low-rank updates to the model’s weights. This approach enables efficient adaptation while preserving the LLM’s general knowledge.
@@
$W=W_{0}+\Delta W,\quad\text{where}\quad\Delta W=BA$ (9)
@@
During fine-tuning, only the weights of the multi-modal projector and the LoRA modules are updated, while all other parameters of the LLM remain frozen. The LLM is initialized with Qwen2.5-7B-Instruct [42], a model pretrained for instruction-following tasks. By jointly optimizing the LoRA parameters and the multi-modal projector, our model effectively adapts to the characteristics of the M3D-Cap and M3D-VQA datasets.

### v3

@@ 1 Introduction
Medical image analysis plays a crucial role in diagnosing and treating diseases such as cancer, cardiovascular conditions, and neurological disorders. However, existing models are often task-specific, lack adaptability to new tasks, and do not support real-time user interactions. Vision-language models (VLMs), such as Contrastive Language-Image Pre-Training (CLIP) [radford2021learning] and Large Language and Vision Assistant (LLaVA) [liu2023visual], offer greater versatility in dynamic clinical settings by aligning medical images with textual reports. CLIP leverages contrastive learning on image-text pairs to enable zero-shot image classification and image-text retrieval. LLaVA extends CLIP by integrating its visual encoder with a large language model (LLM), facilitating interactive tasks such as report generation and visual question answering (VQA). Despite their success in analyzing 2D medical images, such as chest X-rays [endo2021retrieval] and pathology slides [lu2024multimodal], their application to 3D imaging remains limited.
@@
Several VLMs have attempted to bridge this gap, but each has limitations in the 3D setting. For example, PMC-CLIP [lin2023pmc], trained on large-scale biomedical literature images, is restricted to 2D inputs, leading to poor performance on 3D image understanding. RadFM [wu2023towards] unifies 2D and 3D data but is primarily optimized for text generation tasks, such as VQA, and struggles with broader image-text understanding. More recently, M3D-LaMed [bai2024m3d] was introduced as a generalist multi-modal model for 3D medical image analysis. By combining a CLIP-pretrained 3D visual encoder with a 3D spatial pooling perceiver, M3D-LaMed enables direct reasoning on 3D scans and achieves state-of-the-art performance across multiple benchmarks, including image-text retrieval, report generation, open-ended and closed-ended VQA, as well as 3D segmentation and localization tasks.
@@
    Efficient 3D Feature Encoding. We integrate DCFormer [ates2025dcformer], a 3D image encoder that efficiently captures volumetric features by decomposing 3D convolutions into three parallel 1D convolutions along the depth, height, and width axes. This approach reduces computational complexity and enables richer, more scalable feature representations for 3D image volumes.
@@
    Improved Vision-Language Alignment. We adopt SigLIP [zhai2023sigmoid], a sigmoid-based language-image pretraining scheme, to improve image-text alignment. Unlike CLIP’s softmax-based contrastive loss, which relies on distinguishing positive pairs from a large batch of negative samples, SigLIP uses a pairwise sigmoid loss that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, making training more stable and less sensitive to batch size.
@@
    Multi-Scale Multi-Modal Projector. We introduce a novel projector based on MLP-Mixer [tolstikhin2021mlp] to fuse image and text embeddings effectively. Utilizing a low-high-hybrid design, it blends detailed low-level and abstract high-level features from the image encoder with LLM text embeddings using stacked MLP layers. Inspired by MLP-Mixer’s ability to mix spatial and feature data, this dual-stream approach captures richer cross-modal interactions than simple linear projection, improving the LLM’s decoding accuracy.
@@ 2.1 Medical Vision-Language Models
Medical VLMs utilize multi-modal learning to enhance tasks such as disease diagnosis, report generation, and medical VQA through improved integration and understanding of medical images and text. One of the earliest efforts, ConVIRT [zhang2022contrastive], applied contrastive learning to learn visual representations from paried image-text data, outperforming ImageNet pretraining in medical image classification and zero-shot retrieval tasks. Similarly, BioViL [boecking2022making] leveraged large-scale biomedical datasets to refine multi-modal representations, significantly outperforming previous supervised methods. Further extending these capabilities, MedCLIP [wang2022medclip] replaced the InfoNCE loss with semantic matching loss based on medical knowledge, demonstrating its ability to learn generalized representations with limited data.
Recent efforts have aimed to develop VLMs for 3D medical imaging. RadFM [wu2023towards] introduced a generalist VLM trained on massive datasets that consisted of both 2D and 3D medical images and associated radiology reports. It integrated contrastive learning and generative modeling across diverse imaging modalities (X-ray, CT, MRI) and textual reports, enabling a unified radiology representation. Extending this, M3D [bai2024m3d] introduced a large-scale dataset, M3D-Data, along with M3D-LaMed, a multi-modal vision-language model for 3D medical image analysis tasks such as image-text retrieval, report generation, VQA, and promotable segmentation. Its successor, E3D-GPT [lai2024e3d], employed a 3D multi-modal masked autoencoder framework to further enhance image-text retrieval, leading to improved performance in report generation and VQA. In parallel, researchers have also explored text-guided 3D medical image segmentation [zhao2023one, huang2024cat, xin2025text3dsam], where language prompts are used to localize anatomical structures or pathologies.
@@ 2.2 Radiology Report Generation
Automated report generation aims to produce descriptive, accurate, and clinically relevant reports from medical images. It improves diagnostic accuracy while reducing radiologists’ workload and supporting care in resource-limited settings. Early approaches primarily adopted an encoder-decoder architecture, where convolutional neural networks (CNNs) extracted image features, and long short-term memory (LSTM) networks generated text descriptions [xue2018multimodal, harzig2019addressing, yuan2019automatic, wang2021self]. However, these methods struggled with long-range dependencies, often produced repetitive text, and had limited capacity to capture complex medical semantics.
The transformer architecture [vaswani2017attention] addressed key limitations of CNNs and LSTMs by using self-attention to capture global dependencies. Pretrained LLMs [touvron2023llama, penedo2023refinedweb] based on transformers are now adapted into medical VLMs for report generation. R2GenGPT [wang2023r2gengpt] introduced an efficient visual alignment module to better integrate image features with LLM word embeddings, improving text coherence and clinical relevance. Med-Flamingo [moor2023med], based on OpenFlamingo-9B, was pretrained on interleaved medical image-report pairs, achieving superior performance in generating clinically useful responses. More recent methods, such as LLaVA-Med, were based on the LLaVA [liu2023visual] framework, which combined a visual encoder with an LLM to generate detailed and coherent reports. CT2Rep [hamamci2024ct2rep] proposed an advanced 3D vision encoder to generate radiology reports for 3D medical imaging, specifically targeting chest CT volumes. Similarly, CT-CHAT [hamamci2024developing] adapted the LLaVA framework for chest CT report generation, demonstrating the effectiveness of large-scale pretrained LLMs in 3D medical imaging.
@@ 2.3 Medical Visual Question Answering
A medical VQA system can answer natural language questions about medical images. This task can be closed-ended (with answers like yes/no or a choice from a fixed list) or open-ended (free-form text answers). Early work on medical VQA largely treated it as a classification problem, where models selected the correct answer from a predefined set of possible responses [nguyen2019overcoming, zheng2020learning, gupta2021hierarchical, liu2021contrastive]. These models typically encoded images (e.g., using CNNs) and text (e.g., using transformers), and then mapped the combined features to a predefined answer space. Although this approach worked well for simple questions, such as identifying an organ, it struggled with questions requiring detailed explanation or answers not included in the predefined list.
Nowadays, approaches to medical VQA increasingly adopt generative architectures, treating answer generation as a sequence prediction task, which enables open-ended responses. An early step in this direction was CGMVQA [ren2020cgmvqa], which added a generative decoder branch alongside the traditional classifier. Recent VQA systems have fully transitioned to generative models [dong2025generative]. These models (e.g., Med-PaLM M [tu2024towards] and LLaVA-Med [li2023llava]) extend an LLM with visual inputs via visual instruction tuning [liu2023visual], allowing the LLM to interpret medical images and answer questions in a conversational manner. Recently, a few works have explored VQA on 3D medical images [bai2024m3d, hamamci2024developing], where models can directly perceive entire 3D image volumes and generate context-aware responses, offering more accurate and holistic medical insights.
@@ 3.1 Dataset
This study used the publicly available M3D dataset [bai2024m3d], collected with informed consent and ethical approval by the original investigators. All data were de-identified, and no additional consent or approval was required for this secondary analysis. We used two subsets of the M3D dataset: M3D-Cap and M3D-VQA. M3D-Cap includes 120K image-text pairs, while M3D-VQA contains 662K instruction-response pairs. The dataset can support tasks such as image-text retrieval, radiology report generation, and VQA. All image volumes were resampled to a fixed size of 128x256×256. The dataset was divided into training (115k samples), validation (3k samples), and test (2k samples) sets.
@@ 3.2.1 Contrastive Pretraining
Following the CLIP framework [radford2021learning], the vision and text encoders are trained jointly using a contrastive loss to align 3D images with their corresponding radiology reports. The goal is to maximize the similarity between embeddings of positive (matching) image-text pairs while minimizing it for negative (non-matching) pairs.
For the vision encoder, we utilize DCFormer [ates2025dcformer], which efficiently extracts volumetric features using decomposed 3D convolutions. This approach mitigates computational challenges in high-resolution 3D image analysis, where traditional 3D CNNs and vision transformers struggle due to the cubic scaling of 3D CNNs with kernel size and the quadratic scaling of vision transformers with image size. DCFormer decomposes 3D convolutions into three 1D parallel depthwise convolutions:
@@
Such a decomposition strategy enables the processing of 3D volumes at a larger size of 128x256×256 in this study, preserving fine-grained spatial details for improved image-text alignment. DCFormer, short for DeComposed Former, derives its name from this decomposition strategy. The “Former” suffix is inspired by the MetaFormer [yu2022metaformer] framework, which generalizes Transformer-like architectures by replacing attention mechanisms with alternative token-mixing operations.
In this work, we utilize DCFormer-small variant, which consists of a stem stage and four hiararchical stages with [2, 3, 6, 2] layers and [96, 192, 384, 768] channels [ates2025dcformer]. This adaptation of DCFormer within the LLaVA [liu2023visual] framework enables efficient processing of high-resolution 3D volumes while preserving spatial detail. By reducing computational complexity, it supports scalable 3D medical vision-language modeling under typical GPU constraints and addresses key limitations of 3D vision transformers and CNNs.
For the text encoder, we use ClinicalBERT [wang2023optimized], an adaptation of BERT [devlin2019bert] pretrained on clinical notes. Designed for clinical text, ClinicalBERT is well-suited for processing radiology reports, generating embeddings that effectively capture their semantic content.
Traditional softmax-based contrastive learning (e.g., CLIP) distinguishes positive pairs from a large batch of negative samples. While effective for large datasets, this approach struggles with smaller medical datasets, where meaningful negative samples are limited, and similar images may share semantic information. To address this, we adopt SigLIP [zhai2023sigmoid], a sigmoid-based language-image pretraining scheme that directly optimizes each image-text pair independently. This eliminates the need for global similarity normalization across batches, resulting in more stable training and reduced sensitivity to batch size. Sigmoid loss is defined as:
$\mathcal{L}_{\text{SigLIP}}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\sum_{j=1}^{|\mathcal{B}|}\log\underbrace{\frac{1}{1+e^{z_{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}_{\mathcal{L}_{ij}},$ (5)
where $z_{ij}$ is the label for the $i$-th image and $j$-th text pair, $z_{ij}=1$ for positive pairs and $z_{ij}=-1$ otherwise. The sigmoid loss $\mathcal{L}_{ij}$ is computed for each image-text pair, with $t$ as the temperature parameter, $\mathbf{x}_{i}$ and $\mathbf{y}_{j}$ as the image and text embeddings, respectively, and $b$ as a bias term. The overall loss is averaged over all pairs in the batch.
@@ 3.2.2 Multi-modal Projector Pretraining
We propose an effective multi-modal projector for the interaction between visual and textual embeddings. In this pretraining stage, all model weights are frozen except for those in the projector. To effectively capture spatial and feature dependencies, we adopt the MLP-Mixer architecture [tolstikhin2021mlp], which alternates between a token-mixing MLP (spatial mixer) and a channel-mixing MLP.
@@
To further enhance multi-modal fusion, we propose a low-high hybrid Mixer-MLP architecture (Figure 2), inspired by Janus Pro [chen2025janus]. Unlike M3D-LaMed [bai2024m3d], which produces large spatial outputs $(B,2048,768)$ that require lossy downsampling, DCFormer outputs semantically rich features $(B,32,768)$ from its final layer and spatially detailed features $(B,256,384)$ from its penultimate layer. Because of the compact spatial dimensions of DCFormer’s outputs, our model can adopt the low-high hybrid structure without exceeding the token length limit of the LLM. This is crucial, as excessively long image token sequences can crowd out text tokens, leading to truncated questions or reports during inference. To leverage both types of features, we utilize two parallel MLP-Mixer modules—one for each layer. After passing through $N$ Mixer layers, the resulting feature outputs are concatenated into a unified sequence of image tokens. In parallel, the input text is tokenized into a sequence of token IDs using the vocabulary of the LLM. This step follows the LLaVA [liu2023visual] framework, which requires token embeddings for multimodal fusion. The resulting image and text tokens are then fused and passed to the LLM for joint reasoning.
@@ 3.2.3 VLM Fine-tuning
In this stage, we fine-tune the multi-modal projector and the LLM for the report generation and VQA. Instead of fine-tuning the entire LLM, we adopt Low-Rank Adaptation (LoRA) [hu2022lora], which significantly reduces the number of trainable parameters by introducing low-rank updates to the model’s weights. This approach enables efficient adaptation while preserving the LLM’s general knowledge.
@@
$W=W_{0}+\Delta W,\quad\text{where}\quad\Delta W=AB$ (9)
@@
During fine-tuning, only the weights of the multi-modal projector and the LoRA modules are updated, while all other parameters of the LLM remain frozen. The LLM is initialized with Qwen2.5-7B-Instruct [qwen2.5], a model pretrained for instruction-following tasks. By jointly optimizing the LoRA parameters and the multi-modal projector, our model effectively adapts to the characteristics of the M3D-Cap and M3D-VQA datasets.